iosangel:

6m+3= 2m-4/2 Can someone pls guide me through this question.

iosangel:



iosangel:

I know the answer but I forgot the steps...

563blackghost:

what do you think should be done first?

iosangel:

I don't know... Like simplify both sides of the equation?

563blackghost:

well ye, but you gotta chose what to do first. We need to work backwards. Let's first rid of the fraction. Since `2m-4` is being `divided by 2` we will `MULTIPLY by 2` on `both sides`. \(\large\bf{(6m+3) \color{red}{\times 2}=\frac{2m-4}{2} \color{red}{\times 2}}\) This simplifies to what?

iosangel:

Why do we multiply 2 and not subtract?

563blackghost:

There are opposites to each type of sign. When you multiply, the opposite is divide. When you add, the opposite is subtract. When you square root, the opposite is exponent.

iosangel:

Ye I know, but there is 6m+3 so shouldn't we subtract? Like i see the division in the equation....

563blackghost:

We need to work on PEMDAS. So we follow with the fraction first

iosangel:

Ohhhhhhh okay

563blackghost:

cx So what would it simplify to? \(\bf\large{(6m +3) \color{red}{\times 2}=\frac{2m-4}{2} \color{red}{\times 2}}\)

iosangel:

we need to get rid of the fraction, so 6m+3=2m-2?

563blackghost:

Remember that the division of 2 and the multiplying of 2 will cancel the two out.

iosangel:

ye, idk

563blackghost:

You need to apply the multiplication of 2. So... \(\large\bf{(6m \color{red}{\times 2}) + (3 \color{red}{\times 2} )= 2m-4}\)

iosangel:

we distributed here?

563blackghost:

yes

iosangel:

Wait, so (6m+3)*2=2m-4/2*2 simplifies to (6m×2)+(3×2)=2m−4 ?

563blackghost:

well simplify the last part. `6m x 2 = ?` `3 x 2 = ?`

iosangel:

12m and 6

563blackghost:

Correct. So it simplified to: \(\large\bf{12m+6=2m-4}\) Now we need to get the variable to one side (making sure `m` is on one side). Follow by PEMDAS, we see the `addition of 6` to `12` so we want to `SUBTRACT 6` from `BOTH SIDES`. \(\large\bf{12m + 6 \color{red}{-6}=2m-4\color{red}{-6}}\) What does that simplify to?

iosangel:

mmk, lemme copy this down, gemme a sec.

563blackghost:

okie

iosangel:

12m=2m-4-6? Idk

563blackghost:

Okie the left side is correct. Now you just need to do: \(\large\bf{-4-6 ~\rightarrow ~-4 + -6 =?}\)

iosangel:

-10

563blackghost:

ye, so it now becomes: \(\large\bf{12m=2m-10}\) Now we need to get the `variable m` on one side. So we see that `2m` is being subtracted with `10` , BUT `2m` is positive. So we need to `SUBTRACT 2M` from `both sides`. \(\large\bf{12m \color{red}{-2m}=2m \color{red}{-2m} -10}\) What does that simplify to?

iosangel:

10m=-10m

563blackghost:

Well `2m - 2m` equals to 0, so all that would be left is -10.

563blackghost:

So you have: \(\large\bf{10m=-10}\) We need to get the variable to itself. We see that `10` is being `multiplied` to `m`,so we would `DIVIDE` by `10` to get `m` to itself. We do this on `both sides`. \(\huge\bf{\frac{10m}{\color{red}{10}} = \frac{-10}{\color{red}{10}}}\) So what does `m` equal?

iosangel:

-1

563blackghost:

Nice cx

iosangel:

Very much appreciated!!!

563blackghost:

your very welcome cx